OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (1,-3) and has the slope of 2 a) 2x-y=5 b) 2y-y=-5 c)2x+y=5 d) 2x+y=-5

OpenStudy (anonymous):

let y=2x+b it passes through (1,-3) so -3=2*1+b b=-5 y=2x-5 2x-y-5=0 2x-y=5